AB INITIO INVESTIGATION OF THE ELECTRONIC GROUND STATE OF THE NH−N_{2} COMPLEX

\underline{WAFAA M. FAWZY}, Department of Chemistry, East Tennessee State University, Johnson City, TN 37504; MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The NH−N_{2} van der Waals complex has been examined at the CCSD(T) level of theory using the aug-cc-pVDZ basis set. The full basis set superposition error correction was applied. Two minimum energy structures were located for the electronic ground state. The global minimum corresponds to a linear geometry of the complex (NH−N-N), with D_{e}=199 cm^{−1}and R_{cm}=4.3 Å. The secondary minimum corresponds to a T-shaped geometry of C_{2v} symmetry, where the nitrogen atom of the H-N moiety points toward the center of mass of the N_{2} unit, aligned with the a-inertial axis of the complex. The binding energy and R_{cm} value for the secondary minimum were 117 cm^{−1}and 3.7 Å, respectively. Results of the current work on the NH−N_{2} complex will be discussed and compared to results of our previous work on the HN−H_{2} complex^{a}.

aWafaa M. Fawzy, Galina Kerenskaya, and Michael C. Heaven, J. Chem. Phys. 122, 144318, (2005).